You are a member of Board who chairs an ad committee of reforming taxes on telecommunication services. The local telecom tax es can amount to as much as 25 percent of a consumer’s phone bill. The high rates on telecom services have become quite controversial, due to the fact that the deregulation of the telecom industry has led to a highly competitive market. Your best estimates indicate that, based on current tax rates, the monthly market demand for telecommunication services is given by Q=250-5P and the market supply (including taxes) is Q=4P+110 (both in million). The Board of management is considering tax reform that would dramatically cut tax rates, leading to the supply function under the new tax policy of Q=4.171P+110. How much money would typical consumer save each month as a result of proposed legislation?


1
Expert's answer
2021-07-26T10:26:02-0400

Given that:

-The monthly market demand for telecommunication services is given by

"Q_d=250-5p" and the market supply "Q_s=4p+110"

Equilibrium price before tax cut 

"Q_d=Q_s"

"250-5p=4p+110"

"9p=250-110=140"

"9p=140"

"p=\frac {140} {9} =15.55"

p=15.55 per unit.

"Q_d=250-5(15.55)=250-77.75"

Qd=172.25units.

The Supply function under the new tax policy

"Q_s=4.171p+110"

Equilibrium price after tax cut

"Q_d=Q_s"

"250-5p=4.171p+110"

"250-110=4.171p+5p"

"140=9.171p"

"p=\frac {140} {9.171}"

p=$15.26per unit


"Qs=4.171(15.26)+110=63.649+110"

"Q_s" =173.649units


Thus,the money per unit saved by the typical consumer"=15.55-15-26"

=$0.29 per unit